=== 2018/4/19(Thur.) ===

* Remote participants by eZuce: Miyoki

==== Report of work (*:responsible person) ====
 * Phase 1 operation was postponed to 4/28 start and 5/6 end.

 * VAC: *Kimura
  * Large leak was found at the door on the cross tube, and started venting ~15:00. It takes almost one day to start fix the door, will be afternoon today. Pump down will start tomorrow morning.
  * Leak test will be performed on Sunday morning.

 * ETMX: *Hasegawa
  * Transfer function of TM to BF is being investigated.

 * ETMY: *Nakano, Shishido
  * One of the 1.64 and 1.66 Hz peak was suppressed but other one was not. 1-2urad in yaw, 10urad in pitch.
  * RMS motion at REFL port is still dominated by ETMY.

 * BS: *Eleonora, Izumi
  * We are optimizing diagonalization of L2P filter. Needs more time. RMS ~urad


==== Work plan (*:responsible person) ====

 * IFO, ETMs, BS: Enomoto, Nakano, Izumi, Hasegawa, Shishido, Eleonora
  * Damping, BS L2P, restoring alignment.

 * VAC: *Kimura
  * To fix large and small leaks.


==== Laser hazard ====
 * [AM]
  * IR: Upstream SR2-3 line, Xend 1F, Yend 1F
  * Green: PR2, Xend 1F
 * [PM]
  * IR: Upstream SR2-3 line, Xend 1F, Yend 1F
  * Green: PR2, Xend 1F

 * [night]
  * IR: Upstream SR2-3 line, Xend 1F, Yend 1F
  * Green: PR2, Xend 1F


==== Weekly ====

 * week of 4/16
  * Michelson control.
  * Leak check, and fix leaks.

 * week of 4/23
  * Opening the last gate valve.
  * 4/28 Operation start

==== Issues ====